# SHINOBI: Shape and Illumination using Neural Object Decomposition via BRDF Optimization In-the-wild

### 18 Jan 2024 v1

@@ Abstract
Project page: https://shinobi.aengelhardt.com
Figure 1: Object reconstruction using SHINOBI. SHINOBI decomposes challenging in-the-wild image collections into shape, material and illumination using a neural field representation while also optimizing camera parameters.
@@ 3.1 SHINOBI Optimization with Hash Encoding
Camera multiplexes. An effective way to reduce the chance of camera pose optimization to be stuck in local minima is the camera multiplex [14, 27]. For each image, $m$ cameras are jittered around the initial camera and simultaneously optimized. Over time the worst performing camera is repeatedly faded out until $m=1$. This process is visualized in Fig. 3. Since we render multiple proposals for a given image anyway, we see an opportunity to further constrain the optimization using projective geometry. Specifically, we project the point sets $X_{i}$ rendered by the $m-1$ members into the currently highest ranking camera $\Theta_{0}$ of the multiplex using the estimated depth $D_{i}$ from the volumetric rendering. Then we render the projected coordinates using $\Theta_{0}$ and compare the rendered color and alpha values to the ones originally rendered at $\Theta_{1\ldots m-1}$.
@@ 3.2 Losses and Optimization
Optimization. In total, we use three optimizers: One ADAM [36] optimizer for the networks, hash grid embeddings and cameras, respectively. The learning rate is decayed exponentially on all optimizers. In addition to the camera representation and constraints mentioned above the configuration of the camera optimizer turns out to be crucial for success in connection with the hash grids. We use ADAM with the $\beta 1$ value reduced to 0.2 to smooth out the noise in the camera updates. The learning rate is tuned between 1e-3 to 2e-3 depending on scene size. Furthermore, we use three fading $\lambda$ variables to transition the optimization schedule smoothly. Render resolution is continuously increased over the first half of the optimization while the number of active multiplex cameras is reduced. The direct color optimization is faded to the BRDF optimization and the encoding annealing is performed over the first third of the optimization. Focal length updates and the view importance weighting are delayed until an initial shape has been formed. See the supplementary material for a detailed description and visualization of the optimization scheduling.
@@ 4 Experiments
Figure 5: Comparison with SAMURAI decomposition. Due to the improved alignment and representation higher frequency details are reconstructed in shape and the BRDF components compared to SAMURAI. Notice the improved texture detail and silhouettes of ours.
@@
Ablation study. We ablate different aspects of SHINOBI in terms of reconstruction metrics using the “Keywest” and “School Bus”, two in-the-wild sets from NAVI [33] of medium complexity. Metrics in Tab. 3 show that the resolution annealing coarse-to-fine scheme and the patch-based losses contribute most significantly to the final quality. The latter improves local details and registration accuracy compared to a simple pixel-wise loss. The view importance weighting is another important factor for improved sharpness. It helps to stabilize the optimization after the initial resolution annealing schedule has ended. While the hybrid encoding and camera multiplex consistency do not seem to have a large impact quantitatively, they play a critical role in stabilizing the optimization over different scene types and scales. Without them, the optimization might take longer or not converge at all depending on the initialization. Visual examples of the specific ablations are compared in the supplementary material.
Limitations. Joint pose and shape reconstruction is an inherently ill-posed problem. While SHINOBI improves over previous work, especially symmetric objects and highly specular materials can lead to failure cases as shown in Fig. 8. The coarse-to-fine scheme is not able to resolve the disambiguities and the camera poses are stuck in a local minimum. All existing methods show these limitations to some extent. In some regions, high-frequency detail is still not reconstructed properly due to misaligned views and the limited high-frequency capabilities of the illumination representation [13]. Furthermore, our BRDF and illumination decomposition is not capable of modeling shadowing and inter-reflections. As we are mainly concerned with single-object decomposition, the shadows and inter-reflections are not crucial. Extending this method to more complex light transport modeling forms an important future work.
@@ 5 Conclusion
We present SHINOBI, a framework for shape, pose, and illumination estimation of objects from unposed in-the-wild image collections. Our novel hybrid hash grid encoding enables easier camera pose optimization using a multiresolution hash grid. Additionally, our choice of camera parameterization along with per-view importance weighting and patch-based alignment loss allows for a better image-to-3D alignment resulting in better reconstruction with high-frequency details. Although SHINOBI is able to recover the geometry of objects from any category, its performance is limited on thin/transparent structures and fails to recover high-frequency details under extreme illumination changes, which we leave as exploration for future work.
@@ Acknowledgements
This work has been partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645 and SFB 1233, TP 02 - Project number 276693517. It was supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A.
@@ B.2 Additional Visual Results
Fig. 9 shows additional qualitative results on objects from the NAVI dataset compared to the baseline methods. Note, that the methods work at different image resolutions and that we show the original output. NeROIC is able to reconstruct high-frequency detail for scenes that have good initial poses but shows artifacts or fails on others. NeRS suffers from its low resolution mesh representation and often inaccurate camera alignment. SAMURAI and SHINOBI both reproduce appearance that is closer to the original illumination setting due to superior decomposition capabilities while SHINOBI recovers more high-frequency details.
@@ B.3 Qualitative Results of Ablations
Fig. 11 shows qualitative results corresponding to the numerical results from the ablation study reported in the main paper. It can be observed that a robust reconstruction is only possible using the full configuration of our method. While the multiplex consistency loss has only minimal impact on this example the result still shows some visible artifacts and overall increased noise level.

### 29 Mar 2024 v2

@@ Abstract
Figure 1: Object reconstruction using SHINOBI. SHINOBI decomposes challenging in-the-wild image collections into shape, material and illumination using a neural field representation while also optimizing camera parameters. Visit the project page at Project page: https://shinobi.aengelhardt.com
@@ 3.1 SHINOBI Optimization with Hash Encoding
Camera multiplexes. An effective way to reduce the chance of camera pose optimization to be stuck in local minima is the camera multiplex [14, 27]. For each image, $m$ cameras are jittered around the initial camera and simultaneously optimized. Over time the worst performing camera is repeatedly faded out until $m=1$. This process is visualized in Fig. 3. Since we render multiple proposals for a given image anyway, we see an opportunity to further constrain the optimization using projective geometry. Specifically, we project the 2D point sets $X_{i}$ rendered by the $m-1$ members into the currently highest ranking camera $\Theta_{0}$ of the multiplex using the estimated depth $D_{i}$ from the volumetric rendering. Then we render the projected coordinates using $\Theta_{0}$ and compare the rendered color $c_{i}$ and alpha values $\alpha_{i}$ of all cameras in the multiplex to the ones originally rendered at $\Theta_{1\ldots m-1}$.
@@ 3.2 Losses and Optimization
Optimization. In total, we use three optimizers: One ADAM [36] optimizer for the networks, hash grid embeddings and cameras, respectively. The learning rate is decayed exponentially on all optimizers. In addition to the camera representation and constraints mentioned above we use ADAM with the $\beta 1$ value reduced to 0.2 to smooth out the noise in the camera updates. The learning rate is tuned between 1e-3 to 2e-3 depending on scene size. Render resolution is continuously increased over the first half of the optimization while the number of active multiplex cameras is reduced. The direct color optimization is faded to the BRDF optimization and the encoding annealing is performed over the first third of the optimization. Focal length updates and the view importance weighting are delayed until an initial shape has been formed. See the supplementary material for a detailed description and visualization of the optimization scheduling.
@@ 4 Experiments
Figure 5: Comparison with SAMURAI decomposition for joint pose and object reconstruction. Due to the improved alignment and representation higher frequency details are reconstructed in shape and the BRDF components compared to SAMURAI. Notice the improved texture detail and silhouettes of ours. Both methods optimize camera poses jointly initialized from rough quadrants.
@@
Ablation study. We ablate different aspects of SHINOBI in terms of reconstruction metrics using the “Keywest” and “School Bus”, two in-the-wild sets from NAVI [33] of medium complexity. Metrics in Tab. 3 show that the resolution annealing coarse-to-fine scheme and the patch-based losses contribute most significantly to the final quality. The latter improves local details and registration accuracy compared to a simple pixel-wise loss. The view importance weighting is another important factor for improved sharpness. It helps to stabilize the optimization after the initial resolution annealing schedule has ended. While the hybrid encoding and camera multiplex consistency do not seem to have a large impact quantitatively, they play a critical role in stabilizing the optimization over different scene types and scales. Without them, the optimization might take longer or diverge depending on the initialization. Visual examples of the specific ablations are compared in the supplementary material.
@@
Limitations. Joint pose and shape reconstruction is an inherently ill-posed problem. While SHINOBI improves over previous work, especially symmetric objects and highly specular materials can lead to failure cases as shown in Fig. 8. The coarse-to-fine scheme is not able to resolve the disambiguities and the camera poses are stuck in a local minimum. All existing methods show these limitations to some extent. In some regions, high-frequency detail is still not reconstructed properly due to misaligned views and the band limited capabilities of the illumination representation [13]. Furthermore, our BRDF and illumination decomposition is not capable of modeling shadowing and inter-reflections. As we are mainly concerned with single-object decomposition, these are not crucial. Extending this method to more complex light transport modeling forms an important future work.
@@ 5 Conclusion
We present SHINOBI, a framework for shape, pose, and illumination estimation of objects from unposed in-the-wild image collections. Using a hybrid hash grid encoding scheme we enable easier camera pose optimization using a multiresolution hash grid. Additionally, our choice of camera parameterization along with per-view importance weighting and patch-based alignment loss allows for a better image-to-3D alignment resulting in better reconstruction with high-frequency details. Although SHINOBI is able to recover the geometry of objects from any category, its performance is limited on thin/transparent structures and fails to recover high-frequency details under extreme illumination changes, which we leave as exploration for future work.
@@ Acknowledgements
This work has been partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645 and SFB 1233, TP 02 - Project number 276693517.
@@ B.2 Additional Visual Results
Fig. 9 shows additional qualitative results on objects from the NAVI dataset compared to the baseline methods. Note, that the methods work at different image resolutions and that we show the original output. NeROIC is able to reconstruct high-frequency detail for scenes that have good initial poses but shows artifacts or fails on others. NeRS suffers from its low resolution mesh representation and often inaccurate camera alignment. SAMURAI and SHINOBI both reproduce appearance that is closer to the original illumination setting due to superior decomposition capabilities while SHINOBI recovers more high-frequency details. Consequently, on the reference free sharpness metric CPBD [54] our method clearly improves upon SAMURAI and NeRS, with CPBD scores of 0.82 (Ours) vs. 0.77 for SAMURAI vs. 0.65 for NeRS on the most challenging subset of the NAVI scenes where COLMAP reconstruction fails.
@@
Refer to caption (a) Only Fourier Refer to caption (b) Fourier faded out Refer to caption (c) Eval hash grid only Refer to caption (d) Hybrid encoding
Figure 12: Hybrid Encoding Ablation. The Fourier encoding on its own is band limited, e.g. text is not reconstructed. Fading out the Fourier encoding during training destabilizes the optimization showing the importance of both encoding schemes. Evaluation using only the hash grid encoding results in noisy density but sharp texture. Our hybrid encoding yields sharp results with a consistent density at the object’s surface.
@@ B.3 Qualitative Results of Ablations
Fig. 11 shows qualitative results corresponding to the numerical results from the ablation study reported in the main paper. It can be observed that a robust reconstruction is only possible using the full configuration of our method. While the multiplex consistency loss has only minimal impact on this example the result still shows some visible artifacts and overall increased noise level. It is apparent that a plain integration of multi-resolution hash grid encoding does not perform well on the task of joint camera pose and shape reconstruction Fig. 12 visualizes how the encoding schemes complement each other. It has been shown that Fourier encoding together with a large MLP and a coarse-to-fine scheme does perform reasonable well for camera optimization [14, 44]. Hence, we merge the two encoders while keeping the MLP small and therefore band-limited. The main advantage is the continuity of the gradients for the coarse geometry that help to pull the optimization targets closer to the optimum in the beginning of the optimization. Updates based on higher-frequency details later in the optimization can propagate through the hash grids as they are constrained by our additional losses.